OpenStudy (anonymous):

The length of DE is 12. GF = 2x + 2 and FE = 4x - 8. GE is perpendicular to FD. What is the value of x?

OpenStudy (anonymous):

there is an attach file button so that you can take a picture of your problem and upload it. Use these tools.

OpenStudy (anonymous):

(GF)x(FE)=-1 (2x+2)(4x-8)=-1 8x^2 -8x -15=0 Using the quadratic formula we get x=0.25(2-sqrt34) x=0.25(2+sqrt34)

OpenStudy (anonymous):

That can't be right, because the answer choices are 2, 4, 12, and 5. I'll attach a file.

OpenStudy (anonymous):

if that is a parallelogram, then x = 5

OpenStudy (anonymous):

see? it is so much easier to answer questions this way :) I wouldn't have dreamed that it was a quadrilateral. I thought it was a triangle DEF with G being a point on the side DF based on the description you gave.

OpenStudy (anonymous):

Thanks a million, and I'm really sorry for the communication error. Will definitely make sure to be more precise from this point forward.

OpenStudy (anonymous):

That is fine :)